OpenStudy (sissyedgar):

A cruise ship can cover 19 nautical miles in 342 minutes. How many nautical miles will it travel in 234 minutes? A. 17 B. 11 C. 13 D. 15

OpenStudy (sissyedgar):

@sammixboo @texaschic101 @bohotness

OpenStudy (igreen):

Set up a proportion. \(\dfrac{19}{342} = \dfrac{x}{234}\) Multiply 234 to both sides: \(x = \dfrac{19 \times 234}{342}\)

OpenStudy (igreen):

Simplify

OpenStudy (igreen):

Multiply 19 * 234 then divide by 342.

OpenStudy (sissyedgar):

13

OpenStudy (sissyedgar):

ty!

OpenStudy (igreen):

Yep, that's our answer.